# Surgical SAM 2: Real-time Segment Anything in Surgical Video by Efficient Frame Pruning

### v1

@@ 4.3.2 Evaluation on Model Accuracy
Apart from model efficiency, we also evaluated SurgSAM-2’s performance compared with the vanilla SAM2 model. The results, as detailed in Table 1,  2 and 3, indicate that SurgSAM-2 consistently outperforms the vanilla SAM2 across various settings and datasets.
From the experiment results, it is shown that reducing the memory bank size while applying the cosine similarity mechanism led to mixed results in segmentation accuracy for the EndoVis17 dataset. Specifically, while the J&F metric saw a slight decrease of 0.5% and the Dice score dropped by 0.3% compared to the original memory bank configuration, there was an improvement in FPS and memory efficiency. On the other hand, for the more challenging dataset EndoVis18, SurgSAM-2 achieved a 2.2% increase in the J&F metric and a 2.5% improvement in the Dice score, reflecting the positive impact of the proposed EFP mechanism.
@@
However, it’s important to note that while a smaller memory bank can improve performance by reducing noise, it must still be large enough to store all useful information. Striking the right balance is crucial for maintaining a precise and efficient representation of temporal information, ultimately improving overall model performance in surgical video segmentation. The efficiency gains seen with SurgSAM-2 highlight the importance of this balance, as the model achieves improved FPS and reduced memory usage even in cases where segmentation accuracy remains comparable or slightly reduced.
@@ 4.3.3 Fine-Tuning for Optimized Segmentation and Efficiency
We further investigated the efficacy of fine-tuning in our SurgSAM-2 model. Typically, higher input resolutions enhance segmentation accuracy. Surprisingly, we found SurgSAM-2, which predicts segmentation based on half the original resolution, (512, in our setting), can already outperform the vanilla SAM2 at the full resolution. This result is important in the surgical scenario because a smaller resolution allows for real-world model training with reduced memory requirements. Most importantly, this enables a dramatic increase in prediction speed for real-time segmentation of surgical video. In the 1-point setting, which requires very little effort from the surgeon—just a single click on the first frame of the entire surgical procedure—our SurgSAM-2 can increase precision from 85.1% Dice to 87.3%, and also efficiency from 29 to 86 FPS in EndoVis17, as in shown in Table 2.
Looking at all three tables 1,  2 and 3 into details, we can see that our SurgSAM-2 with fine-tuning shows a marked and consistent improvement in segmentation accuracy across different settings. For instance, for Five points, our SurgSAM-2 with fine-tuning led to an increase in the J&F metric from 83.6% to 88.0% and the Dice score from 86.9% to 91.4% in the EndoVis17 dataset. Similarly, our method resulted in an increase in the J&F metric from 76.3% to 80.8% and the Dice score from 80% to 84.9% in the EndoVis18 dataset. This significant enhancement underscores the effectiveness of our fine-tuning strategy in conditions that closely mimic clinical practice, further validating SurgSAM-2’s capability to deliver precise and consistent segmentation in practical medical applications.
@@ 4.3.4 Comparative Model Evaluation
We compared SurgSAM-2 against the state-of-the-art methods specifically designed to surgical instrument segmentation and some advanced SAM-based methods, utilizing the Challenge IoU metric from EndoVis18 to assess performance. Results of other methods are quoted from their papers [14]. Note that a completely fair comparison cannot be achieved, as most of these existing methods do not require prompts in inference. Additionally, most methods for instrument segmentation are designed for type segmentation, which does not need to distinguish different instances of the same type, though this is one of the most challenging problems in surgical instrument segmentation. Our method aims at a more practical setting to segment instruments on an instance level. We also compared the vanilla SAM2 in this setting and listed all the results for the EndoVis18 dataset in Table 4.
@@
SurgSAM-2 (Ours) (1 Point) 72.6
SurgSAM-2 (Ours) (5 Points) 82.1
SurgSAM-2 (Ours) (Full) 84.4
Similar to the performance in other evaluation matrices, we can see SurgSAM-2 can consistently deliver superior results than the vanilla SAM2 in Challenge IoU, whether it is provided with detailed prompts (Full Mask) or more sparse prompts (1 Point and 5 Points). We also find that with the challenging instance-level setting, our method with mask prompt can achieve competitive IoU results, compared with these task-specific methods. More importantly, SurgSAM-2 manages to achieve significant improvements in FPS and memory efficiency, superior to all the other methods.
In real-world surgical environments, both performance and computational demands need to be carefully considered. Given promising segmentation precision, coupled with high efficiency, our SurgSAM-2 shows its great potential to facilitate the applicability of AI models in surgical deployment. Its consistent high performance across different prompting levels makes it a practical choice for integration into medical imaging workflows, providing surgeons with reliable and real-time segmentation results under various conditions.
@@ 4.3.5 Qualitative Evaluation
Fig. 2 presents a qualitative comparison between our SurgSAM-2 and the vanilla SAM2 in full mask, one-point, and five-point prompt settings. The figure demonstrates that the standard SAM2 occasionally fails to segment the target object or identifies the incorrect object. On the contrary, our SurgSAM-2 consistently produces accurate segmentation masks for the target instrument. Incorporating more accurate prompts, like five-point or mask prompts, can further enhance segmentation accuracy.
Figure 2: Visual comparison between SAM2 and the proposed model SurgSAM-2 on EndoVis18 dataset.
@@ 5 Conclusion and Future Work
In conclusion, our proposed model SurgSAM-2 represents a significant advancement in the domain of surgical video segmentation. By integrating an EFP mechanism with the robust SAM2 framework, SurgSAM-2 successfully addresses the challenges of real-time surgical video processing, enhancing both efficiency and accuracy. The ability of SurgSAM-2 to selectively retain the most relevant frames based on cosine similarity has reduced memory usage while simultaneously improving the model’s segmentation performance across various tasks.
Our comprehensive evaluations on the EndoVis17 and EndoVis18 datasets demonstrate that SurgSAM-2 consistently outperforms the vanilla SAM2 model, offering superior processing speed and reduced computational demands without compromising on accuracy. These results suggest that efficient memory management is crucial for advancing video segmentation in resource-constrained environments, particularly in the high-stakes context of surgical interventions.
Looking ahead, future research will focus on refining EFP strategies and experimenting with different memory bank sizes to identify the optimal configuration that maximizes both efficiency and segmentation accuracy. Additionally, we plan to expand the evaluation of SurgSAM-2 across more diverse and complex datasets, further validating its robustness and applicability in various surgical contexts. By continuing to explore and integrate more sophisticated memory management techniques, we aim to push the boundaries of what is possible in real-time video analysis, not only within the medical field but also in broader applications that require rapid and accurate video segmentation.

### v2

@@ 4.3.2 Evaluation on Model Accuracy
Apart from model efficiency, we also evaluated SurgSAM2’s performance compared with the vanilla SAM2 model. The results, as detailed in Table 1,  2 and 3, indicate that SurgSAM2 consistently outperforms the vanilla SAM2 across various settings and datasets.
From the experiment results, it is shown that reducing the memory bank size while applying the cosine similarity mechanism led to mixed results in segmentation accuracy for the EndoVis17 dataset. Specifically, while the J&F metric saw a slight decrease of 0.5% and the Dice score dropped by 0.3% compared to the original memory bank configuration, there was an improvement in FPS and memory efficiency. On the other hand, for the more challenging dataset EndoVis18, SurgSAM2 achieved a 2.2% increase in the J&F metric and a 2.5% improvement in the Dice score, reflecting the positive impact of the proposed EFP mechanism.
@@
However, it is important to note that while a smaller memory bank can improve performance by reducing noise, it should be large enough to store all useful information. Striking the right balance is crucial for maintaining a precise and efficient representation of temporal information, ultimately improving overall performance in surgical video segmentation. The efficiency gains observed with SurgSAM2 underscore the significance of this balance, as the model achieves improved FPS and reduced memory usage without substantially compromising segmentation accuracy.
@@ 4.3.3 Fine-Tuning for Optimized Segmentation and Efficiency
We further investigated the efficacy of fine-tuning in our SurgSAM2 model. Typically, higher input resolutions enhance segmentation accuracy. Surprisingly, we found SurgSAM2, which predicts segmentation based on half the original resolution, (512, in our setting), can already outperform the vanilla SAM2 at the full resolution. This result is important in the surgical scenario because a smaller resolution allows for real-world model training with reduced memory requirements. Most importantly, this enables a dramatic increase in prediction speed for real-time segmentation of surgical video. In the 1-point setting, which requires very little effort from the surgeon—just a single click on the first frame of the entire surgical procedure—our SurgSAM2 can increase precision from 85.1% Dice to 87.3%, and also efficiency from 29 to 86 FPS in EndoVis17, as in shown in Table 2.
Looking at all three tables 1,  2 and 3 into details, we can see that our SurgSAM2 with fine-tuning shows a marked and consistent improvement in segmentation accuracy across different settings. For instance, for Five points, our SurgSAM2 with fine-tuning led to an increase in the J&F metric from 83.6% to 88.0% and the Dice score from 86.9% to 91.4% in the EndoVis17 dataset. Similarly, our method resulted in an increase in the J&F metric from 76.3% to 80.8% and the Dice score from 80% to 84.9% in the EndoVis18 dataset. This significant enhancement underscores the effectiveness of our fine-tuning strategy in conditions that closely mimic clinical practice, further validating SurgSAM2’s capability to deliver precise and consistent segmentation in practical medical applications.
@@ 4.3.4 Comparative Model Evaluation
We compared SurgSAM2 against the state-of-the-art methods specifically designed to surgical instrument segmentation and some advanced SAM-based methods, utilizing the Challenge IoU metric from EndoVis18 to assess performance. Results of other methods are quoted from their papers [27]. Note that a completely fair comparison cannot be achieved, as most of these existing methods do not require prompts in inference. Additionally, most methods for instrument segmentation are designed for type segmentation, which does not need to distinguish different instances of the same type, though this is one of the most challenging problems in surgical instrument segmentation. Our method aims at a more practical setting to segment instruments on an instance level. We also compared the vanilla SAM2 in this setting and listed all the results for the EndoVis18 dataset in Table 4.
@@
SurgSAM2 (Ours) (1 Point) 72.6
SurgSAM2 (Ours) (5 Points) 82.1
SurgSAM2 (Ours) (Full) 84.4
Similar to the performance in other evaluation matrices, we can see SurgSAM2 can consistently deliver superior results than the vanilla SAM2 in Challenge IoU, whether it is provided with detailed prompts (Full Mask) or more sparse prompts (1 Point and 5 Points). We also find that with the challenging instance-level setting, our method with mask prompt can achieve competitive IoU results, compared with these task-specific methods. More importantly, SurgSAM2 manages to achieve significant improvements in FPS and memory efficiency, superior to all the other methods.
In real-world surgical environments, both performance and computational demands need to be carefully considered. Given promising segmentation precision, coupled with high efficiency, our SurgSAM2 shows its great potential to facilitate the applicability of AI models in surgical deployment. Its consistent high performance across different prompting levels makes it a practical choice for integration into medical imaging workflows, providing surgeons with reliable and real-time segmentation results under various conditions.
@@ 4.3.5 Qualitative Evaluation
Fig. 2 presents a qualitative comparison between our SurgSAM2 and the vanilla SAM2 in full mask, one-point, and five-point prompt settings. The figure demonstrates that the vanilla SAM2 occasionally fails to segment the target object or identifies the incorrect object. On the contrary, our SurgSAM2 consistently produces accurate segmentation masks for the target instrument. Incorporating more accurate prompts, such as five-point or mask prompts, can further enhance segmentation accuracy.
Figure 2: Visual comparison between SAM2 and SurgSAM2 on EndoVis18 dataset.
@@ 5 Conclusion and Future Work
In conclusion, our proposed model SurgSAM2 represents a significant advancement in the domain of surgical video segmentation. By integrating an EFP mechanism with the robust SAM2 framework, SurgSAM2 successfully addresses the challenges of real-time surgical video processing, enhancing both efficiency and accuracy. The ability of SurgSAM2 to selectively retain the most relevant frames based on cosine similarity has reduced memory usage while simultaneously improving the model’s segmentation performance across various tasks.
Our comprehensive evaluations on the EndoVis17 and EndoVis18 datasets demonstrate that SurgSAM2 consistently outperforms the vanilla SAM2 model, offering superior processing speed and reduced computational demands without compromising on accuracy. These results suggest that efficient memory management is crucial for advancing video segmentation in resource-constrained environments, particularly in the high-stakes context of surgical interventions.
Looking ahead, future research will focus on refining EFP strategies and experimenting with different memory bank sizes to identify the optimal configuration that maximizes both efficiency and segmentation accuracy. Additionally, we plan to expand the evaluation of SurgSAM2 across more diverse and complex datasets, further validating its robustness and applicability in various surgical contexts. By continuing to explore and integrate more sophisticated memory management techniques, we aim to push the boundaries of what is possible in real-time video analysis, not only within the medical field but also in broader applications that require rapid and accurate video segmentation.